OpenStudy (blueberries1516):

F(-2)=4-2x How did they get answer of 8? Someone wall me step by step please

OpenStudy (photon336):

This is how they got it \[4-2(-2) = 4-(-4) = 8 \]

OpenStudy (blueberries1516):

Oh. Brain fart. Thank you